OpenStudy (anonymous):

It is a common approximation in physics to say that sin x ≈ x near 0. Use a tangent line approximation at 0 to show that this is a good approximation. How much error is there in this approximation at x = π/6? Please help I do not understand this at all

myininaya (myininaya):

So it sounds like we are to find the tangent line to y=sin(x) at x=0

OpenStudy (anonymous):

Yes and then at pi/6 I think

myininaya (myininaya):

Are we suppose to find the tangnet line at pi/6? I don't see that? I see we are to use the approximation we found from the tangent line to approximate sin(pi/6)

myininaya (myininaya):

but maybe I'm wrong. i just don't see where it says to find the tangent line at pi/6

myininaya (myininaya):

anyways which part do you need help on

OpenStudy (anonymous):

At the end it does but it turned into question marks sorry. I need help with all of it

myininaya (myininaya):

so you need help finding the tangent line to y=sin(x) at x=0?

myininaya (myininaya):

well first of all a line looks like y-y1=m(x-x_1) assuming it isn't vertical anyways

myininaya (myininaya):

m=f'(0) (x1,y1)=(0,sin(0))

myininaya (myininaya):

find f' first

OpenStudy (anonymous):

So is f'(0) is sin x=x

myininaya (myininaya):

remember we are trying to find the tangent line to y=sin(x) at x=0 we haven't establish the approximation sin(x)=x yet

myininaya (myininaya):

so we can't use that yet

myininaya (myininaya):

y=sin(x) so y'=?

myininaya (myininaya):

do you know the derivative of sin(x)?

OpenStudy (anonymous):

yes its cos x

myininaya (myininaya):

so y=sin(x) means y'=cos(x) now we just need evaluate y' at x=0 to find the slope at x=0

myininaya (myininaya):

\[y-y_1=f'(0)(x-x_1)\] \[y-y_1=\cos(0)(x-x_1) \\ \text{ and since again we are finding the tangnet line \to f(x)=\sin(x) } \\ \text{ at (0,\sin(0)) then we have the tangnet line is } \\ y-\sin(0)=\cos(0)(x-0)\] I will let you simplify that when you have simplify things you will see how they got the approximation sin(x)=x

OpenStudy (anonymous):

Okay so how do I get to the pi/6

myininaya (myininaya):

you have approximation is sin(x)=x and sin(pi/6) is the exact value pi/6 is the approximation to sin(pi/6) using the sin(x)=x approximation for values near x=0

myininaya (myininaya):

\[relative error=\frac{x_*-x}{x} \] where x* is the approximation to the actual value x

myininaya (myininaya):

you can also find the absolute error \[x_*-x\]

myininaya (myininaya):

they are probably just looking for absolute error

OpenStudy (anonymous):

Okay so what do I plug in for that equation? I'm sorry if I sound really dumb

myininaya (myininaya):

sin(pi/6) is the actual value pi/6 is the approximation...

myininaya (myininaya):

I called x the actual and x_* the approximation

myininaya (myininaya):

so I just plug in those numbers pi/6-sin(pi/6)=absolute error

OpenStudy (anonymous):

Okay makes more sense. Thank you :)

myininaya (myininaya):

now I don't see where in the question it says find the tangent line to y=sin(x) at pi/6 you can if you want to but I think you would be doing extra work to me i read this as find the the tangnet line to f(x)=sin(x) at x=0 and use that approximation sin(x)=x for x=pi/6 just to test how good the approximation is for values near x=0